EY018A

Session: Feb 2012 (session 1)
Wavelength: 6 cm

Comment


To give feedback on this experiment, click here

Experiment Feedback

Torun (Kaz Borkowski at Thu Mar 08 12:06:48 2012) Minor failures:
About 10 minutes, since 4:29 UTC, the telescope was inaccessible due to hang-up of the control system. No other known problems.

Svetloe (I.Stepanova at Thu Mar 08 12:44:17 2012) Success:
No scan No0132 - too late on source. No scan No 0234 - failed due to antenna problem.

Badary (Holkina at Thu Mar 08 12:47:12 2012) Success:
Scans 0017-0019,0021,0022,0024,0025, 0027, 00280, 0030, 0031, 0033, 0034, 0036, 0037 0040, 0042-0064, 0237-0244 failed due to antenna problems.

Medicina (A. Maccaferri at Thu Mar 08 13:23:55 2012) Failure:RECE
Probably error in OL frequency setup

Yebes (Pablo de Vicente at Thu Mar 08 14:25:11 2012) Failure: ANT
We have had severe problems with the control of a part of the frontend that has prevented the system working correctly. Valid data only from 08:49:20. From 07:42 to 08:49 the system has almost worked fine. Tsys was wrong in some parts.

Urumqi (Hua Zhang at Thu Mar 08 19:10:41 2012) Success:

Westerbork (Geert Kuper at Thu Mar 08 20:38:13 2012) Minor failures:
During this whole project we had a mark5B formatter sync error. This will probably result in a large clock offset (139 microsec) but data should be ok.

Shanghai (Bo Xia at Mon Mar 12 04:15:09 2012) Success:

Onsala (Michael Lindqvist at Mon Mar 12 08:20:52 2012) Minor failures: WIND
Started observing at around 08:12 UTC. Wind to high.

Effelsberg (Uwe Bach at Mon Mar 12 11:26:43 2012) Success:

Jodrell_Bank (Alastair Gunn at Mon Mar 19 15:20:52 2012) Success:
No known problems.